OpenStudy (anonymous):

Combine: sqrt 90 - sqrt 10

OpenStudy (anonymous):

sqrt90 = sqrt(9*10) = sqrt(9)*sqrt(10) = 3*sqrt(10) ==== 3*sqrt(10) - sqrt(10) = 2sqrt(10) --- (sqrt(40)